# Kengo Nakamura, Masaaki Nishino, and Shuhei Denzumi NTT Communication Science Laboratories, Kyoto, Japan

### v1

@@ 3.3 Consideration for Element Order
To prove this, we consider the following subfamily.
Definition 12.
@@
In other words, $\mathcal{F}|_{Y,Y^{\prime}}$ is the family of sets generated from $\mathcal{F}$ by first extracting the sets containing every element of $Y$, but no element of $Y^{\prime}$, and then eliminating all of the elements of $Y$ from every set.
The main claim of this section is as follows.
Theorem 13.
@@
If this theorem holds, we can show that the resultant families in Section 3.2 all have an exponential ZDD size regardless of the order of elements. This is because the resultant families in Section 3.2 all have a hidden weighted bit function, a permutation function, or its complements as a subfamily and all of them have an exponential ZDD size regardless of the order of elements; a detailed discussion is given later.
Proof of Theorem 13.
If we can show $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=O(Z_{<}(\mathcal{F})f(m))$ for any $Y,Y^{\prime}\subseteq X$ and any order $<$ of elements, Theorem 13 can be proved as follows: Suppose that there is an order $<$ of elements satisfying $Z_{<}(\mathcal{F})=o(g(m)/f(m))$. Then, by the above equation, we have $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=o((g(m)/f(m))\cdot f(m))=o(g(m))$. This contradicts the assumption that $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=\Omega(g(m))$ for any order $<$ of elements.
@@
The former step can be achieved by the intersection operation. Let $\mathcal{G}$ be the family of subsets of $X$ such that $S\in\mathcal{G}$ if and only if $S$ contains all of the elements in $Y$ but does not contain any element in $Y^{\prime}$. In other words, $\mathcal{G}\coloneqq\{S\subseteq X\mid S\cap Y=Y\wedge S\cap Y^{\prime}=\emptyset\}$. Then, $\mathcal{F}\cap\mathcal{G}$ is the desired family. The ZDD representing $\mathcal{G}$ has the following form: for any $x\in Y$, there is only one ZDD node labeled $x$ whose lo-child is $\bot$ while its hi-child is the next-level node, and for any $x\in X\setminus(Y\cup Y^{\prime})$, there is only one ZDD node labeled $x$ whose lo-child and hi-child are both the next-level node. Thus, we have $Z_{<}(\mathcal{G})=O(f(m))$ because the base set $X$ of $\mathcal{F}$ has $O(f(m))$ elements and, for any element $x\in X$, there is at most one node labeled $x$. Finally, $Z_{<}(\mathcal{F}\cap\mathcal{G})=O(Z_{<}(\mathcal{F})f(m))$.
@@
The latter can be achieved by eliminating the nodes labeled $x\in Y$ and replacing the branches heading it. For a node labeled $x\in Y$, its lo-child must be $\bot$, since the ZDD is reduced and every set in $\mathcal{F}\cap\mathcal{G}$ must contain $x$. For this node, we first make all of the arcs heading to it point to its hi-child. Then, we eliminate this node. By performing this operation for every node labeled $x\in Y$, we finally obtain the ZDD of $\mathcal{F}|_{Y,Y^{\prime}}$. Since this operation does not increase the size of ZDD, we have $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=O(Z_{<}(\mathcal{F})f(m))$. ∎
@@
Now we can show that the resultant families in the proof of Section 3.2 have exponential ZDD size regardless of the order of elements. For example, for the join operation, ($(\{X\}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$ for any order of elements of $Y$ (and thus that of $X\cup Y$). Therefore, by Theorem 13, $Z(\mathcal{F}_{m}\sqcup\mathcal{G}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements of $X\cup Y$. Similar discussions hold for the other operations.
  • For the disjoint join and conjoint join operations, ($(\{X\}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$.
@@
  • For the meet operation, we already have $\mathcal{F}_{m}\sqcap\mathcal{G}_{m}=\mathcal{H}_{m}$, and thus $Z(\mathcal{F}_{m}\sqcap\mathcal{G}_{m})=\Omega(2^{m/5}/m)$ for any order of elements.
@@
  • For the delta operation, $(2^{X}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$, and again $Z(\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m})=\Omega(2^{m/5}/\mathrm{poly}(m))$ for any order of elements.
@@
  • For the quotient operation, $Z(2^{Y})=O(m)$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$ for any order of elements, and $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements by Lemma 7. This also holds for $\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}$, since it equals $\mathcal{H}^{\prime}_{m}$.
@@
  • For the remainder operation, since $\mathcal{F}_{m}=\bigcup_{k}(\{\{x_{k}\}\}\sqcup\mathcal{E}^{\prime}_{m,k})$ and $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})=\{\{x_{1}\},\ldots,\{x_{m}\}\}\sqcup\mathcal{H}^{\prime}_{m}$, we have $\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m}=\bigcup_{k}(\{\{x_{k}\}\}\sqcup(\mathcal{E}^{\prime}_{m,k}\setminus\mathcal{H}^{\prime}_{m}))$. Thus, $(\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m})|_{\{x_{1}\},X\setminus\{x_{1}\}}=\mathcal{E}^{\prime}_{m,1}\setminus\mathcal{H}^{\prime}_{m}$. Here, $Z(\mathcal{E}^{\prime}_{m,1})=O(m^{2})$ and $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements, and $Z(\mathcal{E}^{\prime}_{m,1}\setminus\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{4})$ for any order of elements by Lemma 7. Thus, by Theorem 13, $Z(\mathcal{F}\operatorname{\%}\mathcal{G})=\Omega(2^{m/5}/m^{6})$ for any order of elements.
@@
  • For the permit and nonsubset operations, we already have $\mathcal{F}_{m}\oslash\mathcal{G}_{m}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$ and $\mathcal{F}_{m}\operatorname{\nearrow}\mathcal{G}_{m}=\mathcal{P}_{m}$. Since $Z(\mathcal{C}_{m})=O(m^{3})$ and $\mathcal{P}_{m}=\Omega(2^{m}/m)$ for any order of elements, $Z(\mathcal{C}_{m}\setminus\mathcal{P}_{m})=\Omega(2^{m}/m^{4})$ for any order of elements.
  • For the restrict and nonsuperset operations, we have ($(\{X\}\sqcup(\mathcal{C}_{m}\setminus\mathcal{P}_{m}))|_{X,\emptyset}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$ and ($(\{X\}\sqcup\mathcal{P}_{m})|_{X,\emptyset}=\mathcal{P}_{m}$.
  • For the maximal operation, we have $\mathcal{F}_{m}^{\uparrow}|_{\emptyset,\{w\}\cup X}=\mathcal{P}_{m}$.
  • For the minimal operation, we have $\mathcal{F}_{m}^{\downarrow}|_{\{w\}\cup X,\emptyset}=\mathcal{P}_{m}$.
  • For the minimal hitting set operation, we already have $\mathcal{F}_{m}^{\sharp}=\mathcal{P}_{m}$.
4 Conclusion
We proved that the worst-case complexity of carrying out certain kinds of a family algebra operation on ZDDs once is lower bounded by an exponential factor. These include all of the operations in Table 1 except for the basic set operations, i.e., union, intersection, and difference. In particular, we resolved the controversy over the complexity of the join operation, which had arisen prominently in past literature. Additionally, we also resolved the open problem regarding the worst-case complexity of the quotient operation raised by Knuth [6].
Future directions can be summarized as follows. First, the permutation function is not such a “devilish” example in that it is a family of subsets of a set with $O(m^{2})$ elements and its ZDD size can only be lower bounded by $\Omega(2^{m}/\mathrm{poly}(m))$. We should investigate whether there is a family of sets generated by restrict or similar operations whose ZDD size is lower bounded by $\Omega(\alpha^{n}/\mathrm{poly}(n))$, where $\alpha>1$ and $n$ is the number of elements in the base set. Second, it is unknown whether a “double recursion” procedure like that in Section 2.3 always leads to an exponential worst-case complexity. It is important to investigate whether there are non-trivial operations that should require a double recursion procedure even though the worst-case complexity is polynomially bounded.
Acknowledgments
We thank Hiromi Emoto and Sho Ooba for pointing out the issues regarding the complexity of performing family algebra operations on ZDDs. We also thank Shin-ichi Minato and Jun Kawahara for valuable discussions on this topic. This work was supported by JSPS KAKENHI Grant Number JP20H05963 and JST CREST Grant Number JPMJCR22D3.
Appendix: Polynomially Bounded ZDDs
@@ Acknowledgments
Here, we prove that the ZDD sizes of some families appearing in Section 3 are bounded by a polynomial of $m$. To prove the size bound, we consider the following linear network model to distinguish whether a set is contained in the family $\mathcal{F}$. Note that the idea of a linear network model comes from Knuth’s book [6], where it was used to prove the bound of BDD size. Suppose that the order of elements is $x_{1}<x_{2}<\cdots<x_{n}$. There are $n$ computational modules $M_{1},\ldots,M_{n}$. Module $M_{i}$ receives an input of one bit indicating whether $x_{i}$ is included in the set. Module $M_{i}$ sends $a_{i+1}$ bits of information to module $M_{i+1}$. Overall, every module $M_{i}$ receives an input $x_{i}$ and $a_{i}$ bits of information from $M_{i-1}$ and sends $a_{i+1}$ bits of information to $M_{i+1}$. Since module $M_{1}$ has no preceding module, we set $a_{1}=0$. The final module, $M_{n}$, outputs one bit indicating whether the set is included in the family $\mathcal{F}$. An overview of the linear network model is drawn in Figure 6. The following lemma suggests that if we can construct a small linear network for the family $\mathcal{F}$, the ZDD size of $\mathcal{F}$ can be bounded.
Lemma 14.
@@
We observe that the input to module $M_{i}$ is $a_{i}$ bits. This means that, regardless of the inclusion of $x_{1},\ldots,x_{k-1}$, the subfamily $\mathcal{F}|_{X,Y}$ is completely determined by the information of $a_{i}$ bits. Therefore, there are at most $2^{a_{i}}$ distinct subfamilies, yielding the result that the number of nodes labeled $x_{k}$ is upper-bounded by $2^{a_{i}}$. Since there are two terminal nodes $\top$ and $\bot$, the overall ZDD size is bounded by $Z(\mathcal{F})\leq 2+\sum_{i=1}^{n}2^{a_{i}}$. ∎
@@
By Lemma 14, we only have to consider a small linear network for every family:
$\mathcal{E}_{m,k}$ in Section 3.1.
The family $\mathcal{E}_{m,k}$ is defined as $\{S\subseteq\{y_{1},\ldots,y_{m}\}\mid|S|=k,y_{k}\in S\}$. In judging whether $S\in\mathcal{E}_{m,k}$ with a linear network, the module $M_{t}$ is only concerned with the number of elements from $y_{1},\ldots,y_{t}$ in $S$ and whether $y_{k}$ is in $S$. The former information can be represented with $\lceil\log(m+1)\rceil$ bits and the latter can be represented with 1 bit. Thus, we can construct a linear network with $a_{t}=\lceil\log(m+1)\rceil+1$ bits. By Lemma 14, we have $Z(\mathcal{E}_{m,k})\leq 2+m2^{\lceil\log(m+1)\rceil+1}=O(m^{2})$.
$\mathcal{Q}_{m,k}$ in Section 3.1.
The family $\mathcal{Q}_{m,k}$ $(k=1,\ldots,2m)$ is the family of subsets of $\{y_{1},\ldots,y_{m^{2}}\}$ such that there is exactly one element from a set of $m$ selected elements. In constructing a linear network, the module $M_{t}$ is only concerned with the number of selected elements in $S$: zero, one, or more than one. This information can be represented with $2$ bits. Thus, we have $Z(\mathcal{Q}_{m,k})\leq 2+m^{2}2^{2}=O(m^{2})$.
$\mathcal{E}_{m,k}^{\prime}$ in Section 3.2.2.

### v2

@@ 3.3 Consideration for Element Order
Definition 3.17.
@@
In other words, $\mathcal{F}|_{Y,Y^{\prime}}$ is the family of sets generated from $\mathcal{F}$ by first extracting the sets containing every element of $Y$, but no element of $Y^{\prime}$, and then eliminating all of the elements of $Y$ from every set. This operation is called conditioning and it is a famous result that this can be performed in polynomial time with BDDs [6]. For the sake of completeness, we show this can also be performed in polynomial time with ZDDs, and then we prove the following.
Lemma 3.18.
@@
If this lemma holds, we can show that the resultant families in Section 3.2 all have an exponential ZDD size regardless of the order of elements. This is because the resultant families in Section 3.2 all have a hidden weighted bit function, a permutation function, or its complements as a subfamily and all of them have an exponential ZDD size regardless of the order of elements; a detailed discussion is given later.
Proof 3.19 (Proof of Lemma 3.18).
If we can show $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=O(Z_{<}(\mathcal{F})f(m))$ for any $Y,Y^{\prime}\subseteq X$ and any order $<$ of elements, Lemma 3.18 can be proved as follows: Suppose that there is an order $<$ of elements satisfying $Z_{<}(\mathcal{F})=o(g(m)/f(m))$. Then, by the above equation, we have $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=o((g(m)/f(m))\cdot f(m))=o(g(m))$. This contradicts the assumption that $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=\Omega(g(m))$ for any order $<$ of elements.
@@
The former step can be achieved by the intersection operation. Let $\mathcal{G}$ be the family of subsets of $X$ such that $S\in\mathcal{G}$ if and only if $S$ contains all of the elements in $Y$ but does not contain any element in $Y^{\prime}$. In other words, $\mathcal{G}\coloneqq\{S\subseteq X\mid S\cap Y=Y\wedge S\cap Y^{\prime}=\emptyset\}$. Then, $\mathcal{F}\cap\mathcal{G}$ is the desired family. The ZDD representing $\mathcal{G}$ has the following form: (i) For any $x\in Y$, there is only one ZDD node labeled $x$ whose lo-child is $\bot$ while its hi-child is the next-level node. (ii) For any $x\in Y^{\prime}$, there is no node labeled $x$ by the reduction rule of ZDD. (iii) for any $x\in X\setminus(Y\cup Y^{\prime})$, there is only one ZDD node labeled $x$ whose lo-child and hi-child are both the next-level node. Thus, we have $Z_{<}(\mathcal{G})=O(f(m))$ because the base set $X$ of $\mathcal{F}$ has $O(f(m))$ elements and, for any element $x\in X$, there is at most one node labeled $x$. Finally, $Z_{<}(\mathcal{F}\cap\mathcal{G})=O(Z_{<}(\mathcal{F})f(m))$.
@@
The latter can be achieved by eliminating the nodes labeled $x\in Y$ and replacing the branches heading it. For a node labeled $x\in Y$, its lo-child must be $\bot$, since the ZDD is reduced and every set in $\mathcal{F}\cap\mathcal{G}$ must contain $x$. For this node, we first make all of the arcs heading to it point to its hi-child. Then, we eliminate this node. By performing this operation for every node labeled $x\in Y$, we finally obtain the ZDD of $\mathcal{F}|_{Y,Y^{\prime}}$. Since this operation does not increase the size of ZDD, we have $Z_{<}(\mathcal{F}|_{Y,Y^{\prime}})=O(Z_{<}(\mathcal{F})f(m))$.
@@
Now we can show that the resultant families in the proof of Section 3.2 have exponential ZDD size regardless of the order of elements. For example, for the join operation, ($(\{X\}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$ for any order of elements of $Y$ (and thus that of $X\cup Y$). Therefore, by Lemma 3.18, $Z(\mathcal{F}_{m}\sqcup\mathcal{G}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements of $X\cup Y$. Similar arguments hold for the other operations. We here show that all the resultant families in the proof of Section 3.2 have exponential ZDD size regardless of the order of elements.
Disjoint join $\operatorname{\dot{\bowtie}}$ and joint join $\operatorname{\hat{\bowtie}}$:
The resultant family of these operations in the proof of Theorem 3.5 is ($(\{X\}\sqcup\mathcal{H}_{m})$. Here, ($(\{X\}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$.
Meet $\sqcap$:
@@
In the proof of Theorem 3.5, we already have $\mathcal{F}_{m}\sqcap\mathcal{G}_{m}=\mathcal{H}_{m}$. Thus, $Z(\mathcal{F}_{m}\sqcap\mathcal{G}_{m})=\Omega(2^{m/5}/m)$ for any order of elements.
[…]
@@
In the proof of Theorem 3.5, we have $\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m}=2^{X}\sqcup\mathcal{H}_{m}$. Since $(2^{X}\sqcup\mathcal{H}_{m})|_{X,\emptyset}=\mathcal{H}_{m}$, $Z(\mathcal{F}_{m}\operatorname{\boxplus}\mathcal{G}_{m})=\Omega(2^{m/5}/\mathrm{poly}(m))$ for any order of elements.
[…]
@@
$Z(2^{Y})=O(m)$ and $Z(\mathcal{H}_{m})=\Omega(2^{m/5}/m)$ for any order of elements, and $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements by Lemma 3.7. This also holds for $\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m}$ in the proof of Theorem 3.9 since it equals $\mathcal{H}^{\prime}_{m}$.
[…]
@@
Since $\mathcal{F}_{m}=\bigcup_{k}(\{\{x_{k}\}\}\sqcup\mathcal{E}^{\prime}_{m,k})$ and $\mathcal{G}_{m}\sqcup(\mathcal{F}_{m}\operatorname{/}\mathcal{G}_{m})=\{\{x_{1}\},\ldots,\{x_{m}\}\}\sqcup\mathcal{H}^{\prime}_{m}$, $\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m}=\bigcup_{k}(\{\{x_{k}\}\}\sqcup(\mathcal{E}^{\prime}_{m,k}\setminus\mathcal{H}^{\prime}_{m}))$. Thus, $(\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m})|_{\{x_{1}\},X\setminus\{x_{1}\}}=\mathcal{E}^{\prime}_{m,1}\setminus\mathcal{H}^{\prime}_{m}$. Here, $Z(\mathcal{E}^{\prime}_{m,1})=O(m^{2})$ and $Z(\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{2})$ for any order of elements, and $Z(\mathcal{E}^{\prime}_{m,1}\setminus\mathcal{H}^{\prime}_{m})=\Omega(2^{m/5}/m^{4})$ for any order of elements by Lemma 3.7. Thus, by Lemma 3.18, $Z(\mathcal{F}_{m}\operatorname{\%}\mathcal{G}_{m})=\Omega(2^{m/5}/m^{6})$ for any order of elements because it is a family of subsets of a set with $O(m^{2})$ elements.
[…]
@@
We already have $\mathcal{F}_{m}\oslash\mathcal{G}_{m}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$ and $\mathcal{F}_{m}\operatorname{\nearrow}\mathcal{G}_{m}=\mathcal{P}_{m}$ in the proof of Theorem 3.11. Since $Z(\mathcal{C}_{m})=O(m^{3})$ and $\mathcal{P}_{m}=\Omega(2^{m}/m)$ for any order of elements, $Z(\mathcal{C}_{m}\setminus\mathcal{P}_{m})=\Omega(2^{m}/m^{4})$ for any order of elements by Lemma 3.7.
[…]
We have ($(\{X\}\sqcup(\mathcal{C}_{m}\setminus\mathcal{P}_{m}))|_{X,\emptyset}=\mathcal{C}_{m}\setminus\mathcal{P}_{m}$ and ($(\{X\}\sqcup\mathcal{P}_{m})|_{X,\emptyset}=\mathcal{P}_{m}$; see the proof of Theorem 3.11.
Maximal :
In the proof of Theorem 3.13, we have $\mathcal{F}_{m}^{\uparrow}|_{\emptyset,\{w\}\cup X}=\mathcal{P}_{m}$.
Minimal :
In the proof of Theorem 3.13, we have $\mathcal{F}_{m}^{\downarrow}|_{\{w\}\cup X,\emptyset}=\mathcal{P}_{m}$.
Minimal hitting set :
In the proof of Theorem 3.15, we already have $\mathcal{F}_{m}^{\sharp}=\mathcal{P}_{m}$.
Closure :
In the proof of Theorem 3.15, we have $\mathcal{F}_{m}^{\cap}\cap\mathcal{C}_{m}=\mathcal{P}_{m}$. Since $Z(\mathcal{C}_{m})=O(m^{3})$ for any order of elements, $Z(\mathcal{F}_{m}^{\cap})=\Omega(2^{m}/\mathrm{poly}(m))$ for any order of elements.
3.4 Polynomially Bounded ZDDs
@@
We complete the proof of this section by showing that the ZDD sizes of some families appearing in the previous proofs are bounded by a polynomial of $m$. To prove the size bound, we consider the following linear network model to distinguish whether a set is contained in the family $\mathcal{F}$. Note that the idea of a linear network model comes from Knuth’s book [13, Theorem M], where it was used to prove the bound of BDD size. Suppose that the order of elements is $x_{1}<x_{2}<\cdots<x_{n}$. There are $n$ computational modules $M_{1},\ldots,M_{n}$. Module $M_{i}$ receives an input of one bit indicating whether $x_{i}$ is included in the set. Module $M_{i}$ sends $a_{i+1}$ bits of information to module $M_{i+1}$. Overall, every module $M_{i}$ receives an input $x_{i}$ and $a_{i}$ bits of information from $M_{i-1}$ and sends $a_{i+1}$ bits of information to $M_{i+1}$. Since module $M_{1}$ has no preceding module, we set $a_{1}=0$. The final module, $M_{n}$, outputs one bit indicating whether the set is included in the family $\mathcal{F}$. An overview of the linear network model is drawn in Figure 4. The following lemma suggests that if we can construct a small linear network for the family $\mathcal{F}$, the ZDD size of $\mathcal{F}$ can be bounded.
Lemma 3.20.
@@
We observe that the input to module $M_{i}$ is $a_{i}$ bits. This means that, regardless of the inclusion of $x_{1},\ldots,x_{k-1}$, the subfamily $\mathcal{F}|_{X,Y}$ is completely determined by the information of $a_{i}$ bits. Therefore, there are at most $2^{a_{i}}$ distinct subfamilies, yielding the result that the number of nodes labeled $x_{k}$ is upper-bounded by $2^{a_{i}}$. Since there are two terminal nodes $\top$ and $\bot$, the overall ZDD size is bounded by $Z(\mathcal{F})\leq 2+\sum_{i=1}^{n}2^{a_{i}}$.
@@
By Lemma 3.20, we only have to consider a small linear network for every family.
$\mathcal{E}_{m,k}$ in Section 3.1:
The family $\mathcal{E}_{m,k}$ is defined as $\{S\subseteq\{y_{1},\ldots,y_{m}\}\mid|S|=k,y_{k}\in S\}$. In judging whether $S\in\mathcal{E}_{m,k}$ with a linear network, the module $M_{t}$ is only concerned with the number of elements from $y_{1},\ldots,y_{t}$ in $S$ and whether $y_{k}$ is in $S$. The former information can be represented with $\lceil\log(m+1)\rceil$ bits and the latter can be represented with 1 bit. Thus, we can construct a linear network with $a_{t}=\lceil\log(m+1)\rceil+1$ bits. By Lemma 3.20, we have $Z(\mathcal{E}_{m,k})\leq 2+m2^{\lceil\log(m+1)\rceil+1}=O(m^{2})$.
$\mathcal{Q}_{m,k}$ in Section 3.1:
Each of the families $\mathcal{Q}_{m,k}$ $(k=1,\ldots,2m)$ is the family of subsets of $\{y_{1},\ldots,y_{m^{2}}\}$ such that there is exactly one element from a set of $m$ selected elements. In constructing a linear network, the module $M_{t}$ is only concerned with the number of selected elements in $S$: zero, one, or more than one. This information can be represented with $2$ bits. Thus, we have $Z(\mathcal{Q}_{m,k})\leq 2+m^{2}2^{2}=O(m^{2})$.
$\mathcal{E}_{m,k}^{\prime}$ in Section 3.2.2:
@@
3.5 Discussion
Finally, we give some discussions for the presented results. First, we argue theoretical results for BDDs. As stated in Lemma 2.2, the sizes of BDD and ZDD differ only by a linear factor of the size of the base item set. All the results in Section 3.2 have the same form that the number of elements is $O(\mathrm{poly}(m))$, the input ZDD sizes are $O(\mathrm{poly}(m))$, and the output ZDD size is exponential in $m$. Therefore, even if these families are represented by BDDs, the input BDD sizes are all $O(\mathrm{poly}(m))$, and the output BDD sizes are all exponential in $m$. Moreover, the output BDD sizes remain exponential in $m$ for any order $<$ of elements since Lemma 2.2 holds for any order $<$ of elements. This constitutes the theoretical result that the family algebra operations in Table 1, except for the first four operations, cannot be performed in polynomial time in the input BDD sizes.
Second, we discuss how often such exponential blow-up occurs. Although we rely on specific families, the hidden weighted bit function $\mathcal{H}_{m}$ and the permutation function $\mathcal{P}_{m}$, the heart of the above proofs is that even a single operation may cause us to compute the union or intersection of multiple subfamilies. Apart from these families, it is usual that taking the union or intersection of multiple families leads to exponential blow-up. To imagine this, we consider encoding a family described by polynomial-sized conjunctive normal form (CNF) into BDD/ZDD. Every clause can be encoded into a polynomial-sized BDD/ZDD. Moreover, if the entire CNF is encoded into BDD/ZDD, we can solve SAT, or even more difficult #SAT, in linear time with respect to the size of BDD/ZDD [13]. However, it is a famous fact that SAT and #SAT are in NP-complete and #P-complete, respectively, meaning that they are believed not to be solved in polynomial time. This means that for many CNFs, the BDD/ZDD after taking intersection of clauses does not remain polynomial-sized. Therefore, apart from the specific examples used in the proof, there are many cases yielding the blow-up of BDD/ZDD size after single family algebra operation.
Finally, we argue the limitation of some of the above results that the permutation function is not such a “devilish” example. The permutation function is a family of subsets of a set with $O(m^{2})$ elements and its ZDD size can only be lower bounded by $\Omega(2^{m}/\mathrm{poly}(m))$. Since the ZDD size of the family of subsets of a set with $O(m^{2})$ can be at most $\Omega(2^{m^{2}}/\mathrm{poly}(m))$, it is far from being the worst-case. We should investigate whether there is a family of sets generated by restrict or similar operations whose ZDD size is lower bounded by $\Omega(\alpha^{n}/\mathrm{poly}(n))$, where $\alpha>1$ and $n$ is the number of elements in the base set.
4 Conclusion
We proved that the worst-case complexity of carrying out certain kinds of a family algebra operation on BDDs/ZDDs once is lower bounded by an exponential factor. These include all of the operations raised by Knuth [13, §7.1.4 Ex. 203,204,236,243] except for the basic set operations. In particular, we resolved the controversy over the complexity of the join operation, which had arisen prominently in past literature. We also resolved the open problem regarding the worst-case complexity of the quotient operation.
Future directions include the followings. First, we only prove the lower-bound of the complexity of carrying out a single operation. It should be investigated whether we can obtain a non-trivial upper-bound of the complexity. Second, it is unknown whether a “double recursion” procedure like those in Section 2.2 always leads to an exponential worst-case complexity. It is important to investigate whether there are non-trivial operations that should require a double recursion procedure even though the worst-case complexity is polynomial.